Bibliography

compendium

23/03-04 16:19

[doob1993a] Michael Doob. A gentle introduction to TEX. Technical report, Department of Mathematics, The University of Manitoba, 1993.

[eckel1998a] Bruce Eckel. Thinking in Java. Prentice Hall, 1998.

[grimley19??a] Allan Grimley, Steve Hill, Richard Jones, Simon Thomson, and David Turner. Functional programming in Miranda, lecture notes. Computing Laboratory, University of Kent at Canterbury, UK.

[grubert2000a] Nico Grubert. Z Class Tutorial. Beehive GmbH, 2000.

[grune1990a] Dick Grune and Ceriel J.H. Jacobs. Parsing Techniques – A Practical Guide. Ellis Horwood Limited, 1990.

[hartvigsen1990b] Gunnar Hartvigsen, Randi Karlsen, and Tryggve Sørensen. Datafaglig etikk. Innstilling fra etikkutvalget ved seksjon for datafag, Seksjon for Datafag, Institutt for Matematiske Realfag, Universitetet i Tromsø, December 1990.

[hartvigsen1996a] Gunnar Hartvigsen. Forskerhåndboken 1. Universitetet i Tromsø, January 1996. Versjon 5.2.

[hartvigsen1996b] Gunnar Hartvigsen. Forskerhåndboken 2. Universitetet i Tromsø, January 1996. Versjon 5.2.

[jervell1989a] Herman Ruge Jervell. Forelesninger i logikk. Kompendium 35, Institutt for Informatikk, Universitetet i Oslo, April 1989.

[jolitz1991a] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386. Dr. Dobb’s Journal, 1991/1992.

[kirchner2000a] Katrin Kirchner. Zope: Users, Roles and Permissions. Beehive GmbH, 2000.

[kristiansen1995a] Lars Kristiansen. Første ordens teorier og grunnleggende rekursjonsteori. Kompendium 66, Universitetet i Oslo, Institutt for Informatikk, Vårsemesteret 1995.

[manna1992a] Zohar Manna and Amir Pnueli. The Temporal Logic of Reactive and Concurrent Systems. Springer-Verlag, 1992.

[milner1980a] Robin Milner. A Calculus of Communication Systems, volume 92 of Lecture Notes in Computer Science. Springer-Verlag, 1980.

[ness1992a] Bjørn Ness. Into the Great Wide Open — ABC om kommunikasjonstjenester. USIT, Universitetet i Oslo, UNINETT edition, 1992.

[oetiker1999a] Tobias Oetiker, Hubert Partl, Irene Hyna, and Elisabeth Schlegl. The Not So Short Introduction to L^{A}TEX2e, April 1999. Version 3.7.

[omg1998a] Object Managment Group. The common object request broker: Architecture and specification. Technical report, Object Managment Group, February 1998. (revision 2.2).

[omg1999a] Object Managment Group. The common object request broker: Architecture and specification. Technical report, Object Managment Group, October 1999. (revision 2.3.1).

[ostroff1989a] Jonathan S. Ostroff. Temporal Logic for Real-Time Systems. Advanced Software Development Series. Research Studies Press/John Wiley & Sons, 1989.

[pierce1991a] Benjamin C. Pierce. Basic Category Theory for Computer Scientists. The MIT Press, 1991.

[poigne1992a] Axel Poigné. Basic category theory. In S. Abramsky, Dov M. Gabbay, and T. S. E. Maibaum, editors, Handbook of Logic in Computer Science, pages 413–637. Clarendon Press, 1992.

[stensby1999a] Åge Stensby. Internett tjenester 3ip80, kompendium. Technical report, Høgskolen i Hedmark, 1999.

[uito1991a] Universitetet i Tromsø. D-102 artikkelsamling. Seksjon for Datafag, Institutt for Matematiske Realfag, Universitetet i Tromsø, 1991.

[uito1995a] Universitetet i Tromsø. Multimedia systemer. Artikkelsamling ODS-seminar Høst 95, 1995.

[uito1996a] Universitetet i Tromsø. V-301 artikler. Samling av artikler fra V-301 våren 1996, 1996.